E13 Tombert Decorating uses a job order cost system to collect the costs of its interior decorating business. Each client’s consultation is treated as a separate job. Overhead is applied to each job based on the number of decorator hours incurred. Listed below are data for the current year.

Estimated overhead $960,000
Actual overhead $982,800
Estimated decorator hours 40,000
Actual decorator hours 40,500

The company uses Operating Overhead in place of Manufacturing Overhead. Instructions

(a) Compute the predetermined overhead rate.
(b) Prepare the entry to apply the overhead for the year.
(c) Determine whether the overhead was under- or overapplied and by how much.
Answer #1
a
Predetermined overhead rate = 960000/40000= $24 per decorator hour
b
Work in process inventory 972000 =40500*24
        Operating Overhead 972000
c
Under applied overhead = 982800-972000 = $10800